### v2

# ISRE-$0$-$0$.00

[…]
ISRE-$0$-$0$.00
[…]
Learning to Adopt Generative AI
[…]
Learning to Adopt Generative AI
[…]
Recent advancements in generative AI, such as ChatGPT, have dramatically transformed how people access information. Despite its powerful capabilities, the benefits it provides may not be equally distributed among individuals—a phenomenon referred to as the digital divide. Building upon prior literature, we propose two forms of digital divide in the generative AI adoption process: (i) the learning divide, capturing individuals’ heterogeneous abilities to update their perceived utility of ChatGPT; and (ii) the utility divide, representing differences in individuals’ actual utility derived from per use of ChatGPT. To evaluate these two divides, we develop a Bayesian learning model that incorporates demographic heterogeneities in both the utility and signal functions. Leveraging a six-month clickstream dataset, we estimate the model and find significant learning and utility divides across various demographic attributes. Interestingly, lower-educated and non-white individuals derive higher utility gains from ChatGPT but learn about its utility at a slower rate. Furthermore, males, younger individuals, and those with an IT background not only derive higher utility per use from ChatGPT but also learn about its utility more rapidly. Besides, we document a phenomenon termed the belief trap, wherein users underestimate ChatGPT’s utility, opt not to use the tool, and thereby lack new experiences to update their perceptions, leading to continued underutilization. Our simulation further demonstrates that the learning divide can significantly affect the probability of falling into the belief trap, another form of the digital divide in adoption outcomes (i.e., outcome divide); however, offering training programs can alleviate the belief trap and mitigate the divide. Our findings contribute to the literature on the digital divide, AI adoption, and Bayesian learning. We also provide practical implications for governments, Non-profit organizations, AI providers, and companies aiming to promote equitable AI adoption.
[…]
Generative AI, AI Adoption, Digital Divide, Bayesian Learning, Human-AI interaction
1 Introduction
1.1 Motivation
Search engines have long been the primary means by which individuals access online resources; over 93% of internet activities start with a search query [^1] See https://www.webfx.com/seo/statistics/. For instance, Google, which dominates the market with approximately 90% share, processes about 2 trillion searches annually [^2] See https://blog.hubspot.com/marketing/google-search-statistics. However, the recent emergence of large language models (LLMs) offers a new means to access these resources. Trained on extensive datasets (Zhang et al. 2024), applications such as ChatGPT can provide natural language responses to a wide range of user queries (Wang et al. 2023). The general-purpose nature of generative AI powered by LLMs (Morris et al. 2023), combined with their vast knowledge base, positions them as a compelling alternative to search engines (Xu et al. 2023).
Considering the significant potential of generative AI, one might anticipate widespread benefits. However, these potential advantages may be still constrained by the “digital divide,” which refers to the disparities in access, capabilities, and benefits derived from digital tools among individuals of varied backgrounds (Wei et al. 2011). Understanding and addressing this issue is crucial when aiming to transform technological progress into broader economic and societal advancements while ensuring equity and inclusiveness across the population (Burton-Jones and Sarker 2023, Simchi-Levi 2020).
Studying the digital divide is challenging due to both the complexity of the issue and the methodologies employed. Importantly, disparities in digital access and usage manifest differently across various technologies, making demographic differences context-dependent (Haslberger et al. 2023). As technology advances, research into these disparities has evolved from focusing on internet access to exploring various digital platforms and, more recently, artificial intelligence—leading to a new subfield termed the “AI divide” in the context of traditional AI (McElheran et al. 2024). However, given the uniqueness and power of LLM-based generative AI, which represents the first generation of artificial general intelligence (AGI) (Morris et al. 2023), it is imperative to understand the associated disparities and potentially initiate further research in the context of generative AI.
1.2 Research Questions and Findings
Building on previous research on the digital divide, we also conceptualize the GenAI divide in three layers (Wei et al. 2011, Dewan and Riggins 2005): (1) the access divide, referring to inequalities in technology access; (2) the capability divide, concerning individual differences in ability to effectively utilize the tools once accessed; and (3) the outcome divide, describing disparities in final outcomes, such as productivity and performance, which the preceding layers and other factors may influence. Furthermore, literature on IT and AI adoption has long emphasized the importance of individuals’ perceptions during the adoption process (e.g., perceived usefulness and perceived ease of use), as perceived benefits influence adoption decisions (Jo and Bang 2023, Venkatesh et al. 2003). However, perceptions are dynamic and inherently intangible, making it difficult to monitor how and why they change using data from surveys or experiments. Given these methodological challenges, it is understandable that previous digital divide research has not addressed disparities in individuals’ ability to effectively update their perceived utility through repeated interactions and other informational resources. We term this disparity the “learning divide,” a novel aspect of the second-layer capability divide. The learning divide is vital because varying abilities to update beliefs influence actual adoption and usage decisions, thereby affecting the welfare individuals ultimately receive (i.e., outcome divide). For example, if an individual erroneously believes the system is completely useless after a single trial, they may cease using it, thereby forfeiting opportunities to enhance their work and quality of life through the use of new technology. Conversely, if another person forms an accurate understanding after initial use, they can employ the system effectively and maximize their gains. To empirically investigate the existence of our proposed learning divide, we pose the following research question:
RQ1 (learning divide): How do individuals with different demographic backgrounds update their perceived utility of generative AI at varying rates?
Given that individuals’ perceptions are inherently unobservable and must be inferred from data, and their adoption decision-making is subject to informational uncertainty, structural learning models are ideally suited to address this problem (Wang and Yang 2024, Fang 2022). Therefore, to address the first research question, we develop a customized structural learning model to characterize individuals’ dynamic learning and adoption processes of generative AI, allowing for heterogeneous learning speeds related to their demographic attributes.
In addition to the newly defined learning divide, individuals’ disparate utility derived from generative AI upon usage has become the focus of recent studies, although findings are mixed. On one hand, earlier research on AI’s impact on productivity suggests that workers with less prior expertise gain greater benefits from utilizing ChatGPT for specific tasks, potentially ameliorating existing inequalities (Noy and Zhang 2023). On the other hand, studies on AI adoption emphasize the significance of AI literacy (Ng et al. 2021), and note that performance disparities between older and younger groups have widened with the adoption of generative AI like ChatGPT (Haslberger et al. 2023). Most of these works are conducted in controlled experimental settings where participants are required to use specific tools to complete predetermined tasks. However, it remains underexplored how individuals perform in real-world contexts when they can potentially use general-purpose technologies for diverse tasks. To supplement previous studies, we aim to uncover the differences in utility gains when individuals from various backgrounds use generative AI tools. We term this the “utility divide,”, and pose the following research question:
RQ2 (utility divide): How do individuals with different demographic backgrounds gain utility differently from generative AI when they opt to use it?
Consistent with prior digital divide research employing behavioral experiments (Kohli et al. 2024, Haslberger et al. 2023, Robinson et al. 2020), our utility divide evaluates disparities in utility per use, conditional on adoption, constituting a specific type of third-level outcome divide. However, investigating this phenomenon in a real-world context is inherently challenging due to the natural feedback loop in the adoption process. When individuals believe a technology has high utility, they are more likely to adopt it. Upon using the generative AI tool, users update their beliefs based on their experiences (signals within the structural learning model), which in turn influence subsequent adoption decisions. Since individuals lack a priori knowledge of the actual utility, it influences their adoption decisions only indirectly through signal realizations, which are also governed by their learning abilities. Therefore, recognizing that the utility divide and learning divide are intertwined, we additionally incorporate utility heterogeneity into the structural learning framework to jointly model both disparities.
To empirically estimate our model and address the abovementioned research questions, we obtained a dataset from a leading data collection firm. The dataset comprises online footprints of thousands of panelists over a six-month period following the launch of ChatGPT, enabling us to observe individual usage patterns over time. Additionally, the company provided demographic attributes for each participant, allowing us to explore heterogeneity. After fitting our model to the dataset, we uncovered significant and noteworthy learning and utility divides. Specifically, we find that male, younger, and individuals with an IT background both derive higher utility per use from ChatGPT and learn about its utility more rapidly. Besides, although individuals from minority racial groups or with lower educational attainment tend to derive greater utility gains per use, they unfortunately update their beliefs more slowly.
Although we have empirically verified both the learning divide and the utility divide, it remains unclear how these disparities influence individuals’ lifetime usage and welfare dynamics—another form of outcome divides. The utility divide is, by definition, straightforward to interpret since it directly measures the utility gain per usage. However, the impact of the learning divide is less evident, as it requires translating differences in learning speed into variations in adoption dynamics. Furthermore, it remains unclear how effective counterfactual policies, such as training programs, might be in influencing these outcome divides through the learning process. Addressing this gap also aligns with the call for actionable insights in the digital divide literature (Wang et al. 2024, Wei et al. 2011). Therefore, we propose our final research question:
RQ3: How does the learning divide affect the outcome divide, and how to mitigate this outcome disparity?
To address this research question, we leverage our estimated model parameters to simulate individuals’ belief evolution trajectories over time. Through these simulations, we identify a phenomenon we term the “belief trap” wherein users who develop initially low perceptions of generative AI’s utility may cease usage, thereby forgoing opportunities to update their beliefs through further interactions. This self-reinforcing cycle bears striking similarities to the “poverty trap” in development economics, where individuals or nations become locked in persistent poverty without external intervention (Banerjee 2011). Our analysis reveals that users with disadvantaged learning abilities—as identified through our structural model—face significantly higher probabilities of falling into this belief trap, representing another manifestation of the outcome divide. Furthermore, our counterfactual simulations demonstrate that targeted training programs aimed at increasing exposure to generative AI can effectively mitigate disparities in trapping probabilities, indicating a viable policy intervention to address this form of the digital divide.
1.3 Contributions
This study makes several theoretical contributions. First, we present a unified framework for understanding IT adoption dynamics and the digital divide, particularly in scenarios characterized by information uncertainty and learning dynamics. This model is extensible to the adoption of other emerging technologies beyond generative AI. Second, we introduce two novel theoretical constructs—the learning divide and the belief trap—which are broadly applicable and offer new perspectives on technology adoption barriers. Third, within the emerging literature on generative AI, we are among the first to document demographic-based disparities in utility per use and learning abilities among individuals, quantifying how these differences manifest in adoption outcomes through our structural model.
Our paper also generates practical implications for multiple stakeholders. For generative AI providers, the presence of the utility gap underscores the imperative to implement inclusive design practices to enhance the tool’s value for underprivileged users. Moreover, our observation of learning disparity and belief trap suggests the importance of reducing information asymmetry through tutorials and training programs. This responsibility also extends to governments and non-profit organizations (NGOs). Those institutions can leverage our findings to design training programs and outreach initiatives that specifically target demographic groups we identified as susceptible to slower learning or belief traps.
1.4 Organization of the Paper
The remainder of the paper is organized as follows. Section 2 offers a comprehensive literature review and situates our work within various research streams. Section 3 describes the data sources, variable construction, and exploratory analyses, underscoring the necessity of a structural learning framework. In Section 4, we present the model in detail, with estimation procedures and identification strategies outlined in Section 5. Section 6 provides the model estimation results and addresses RQ1 and RQ2. In Section 7, we perform further numerical experiments and evaluate counterfactual policy scenarios in response to RQ3. Finally, Section 8 concludes with broader discussions and highlights future research directions.
2 Related Literature
This research contributes to three streams of literature: the digital divide, AI adoption, and dynamic learning processes. In the subsequent subsections, we review findings from previous literature and highlight our study’s novel additions.
2.1 Digital Divide
The digital divide refers to disparities in access to, capability with, and effective use of information and communication technologies (ICT) among individuals from diverse backgrounds (Vassilakopoulou and Hustad 2023, OECD 2001). As a critical societal issue, it has been extensively researched (see reviews by Vassilakopoulou and Hustad (2023), Lythreatis et al. (2022), Scheerder et al. (2017), Riggins and Dewan (2005)), with key references summarized in Appendix Table LABEL:tab:digitaldivide. Owing to its complex, multidimensional nature, the digital divide manifests at three distinct levels (Wei et al. 2011, Dewan and Riggins 2005): (1) the first-level digital access divide, concerning inequalities in access to ICT; (2) the second-level digital capability divide, addressing disparities in the ability to effectively use ICT among those with access; and (3) the third-level digital outcome divide, describing differences in outcomes due to variations in access, usage patterns, and other related factors (Wei et al. 2011).
The literature has identified a series of key factors contributing to the digital divide (Lythreatis et al. 2022, Van Dijk and Hacker 2003). These factors operate across multiple levels and dimensions, including socio-demographic and socio-economic characteristics (e.g., gender and age) (Venkatesh et al. 2014, Fairlie 2004, Jackson et al. 2001); personal attributes and skills (e.g., motivation and personality traits) (Celik 2023, Hsieh et al. 2011, Hargittai 2002); social and environmental influences (e.g., social support networks and geographical location) (Kohli et al. 2024, Hargittai 2003); and technical factors (e.g., platform affordance) (Wang et al. 2024, Yang and Zhang 2023). Among these, socio-demographic and socio-economic characteristics have consistently emerged as significant determinants of the digital divide (Venkatesh et al. 2014, Fairlie 2004, Jackson et al. 2001). Factors such as education level (Hidalgo et al. 2020, Hsieh et al. 2011, 2008), income (Fairlie 2004), age (Jackson et al. 2001), gender (Van Dijk and Hacker 2003, Hargittai 2003), and ethnicity (Fairlie 2004) have been shown to significantly influence individuals’ access to ICT, their ability to effectively use it, and the benefits they derive from ICT across different contexts and over time (Celik 2023, Venkatesh et al. 2014).
However, it is important to note that the manifestation of the digital divide exhibit substantial variation across technologies and platforms (Yang and Zhang 2023, Burtch and Chan 2019), and that demographic effects on adoption and usage patterns may differ markedly depending on the specific technology or context being examined (Haslberger et al. 2023). As such, the digital divide is inherently a context-specific concept, which makes it necessary to specify the type of ICT under investigation (Vassilakopoulou and Hustad 2023). Consequently, the discourse surrounding the digital divide has evolved alongside the progression of digital technologies. Early studies focus on the unequal access to and use of personal computers (Autor et al. 1998), while later researcher expanded to include Internet access and use (Van Dijk and Hacker 2003, Hargittai 2002). Subsequently, the focus shifted towards the use of online services, such as e-government platforms (Venkatesh et al. 2014), and more recently has encompassed digital platforms including medical crowdfunding (Burtch and Chan 2019) and online rental housing markets (Boeing 2020).
The proliferation of AI technologies has led to a new dimension of digital inequality, the AI divide (McElheran et al. 2024). However, existing literature has primarily examined this divide within the context of narrow, task-specific traditional AI (Celik 2023), leaving a gap in our understanding of disparities in generative AI adoption. Generative AI such as ChatGPT, powered by large language models (LLMs), is distinguished by its foundation in general-purpose models and its ability to produce flexible, natural language outputs. This technology facilitates intuitive human-AI interactions that require minimal technical expertise from users, thereby broadening access for a diverse range of individuals. However, the outputs of generative AI are characterized by inherent variability (Morris et al. 2023). This uncertainty suggests that the utility derived from these systems can vary significantly among users with different backgrounds. Furthermore, effectively evaluating the utility of such AI systems may require an extensive learning process involving repeated interactions. Given the unique characteristics of generative AI, prior findings on the general digital divide or the specific AI divide may not be directly applicable to the context of generative AI.
Our study addresses this gap by developing a unified framework to address both the antecedents and the consequences of AI adoption. Because individuals’ perceived utility from AI influences their adoption decisions, and the realized outcomes subsequently affect their perceptions of AI, a natural feedback loop arises. However, due to the dynamic and intangible nature of perceptions, prior behavioral studies face methodological challenges in directly capturing the utility underlying participants’ adoption decisions or in addressing the dynamic adoption process. To move beyond the “snapshots” provided by survey studies, we develop a Bayesian learning model that offers theoretical underpinnings for the dynamics of the generative AI adoption process, complementing previous behavioral studies. Our model leverages a rich real-world dataset to recover the underlying drivers of adoption decisions (signals and perceived utility) and the direct consequences of individual adoption (actual utility). The model estimates can also be employed for counterfactual simulations to comprehensively evaluate indirect consequences, such as belief traps, and to assess policy interventions. Furthermore, our proposed model can be leveraged for future AI adoption studies involving other factors or technologies.
2.2 AI Adoption
Our paper contributes to the emerging literature on AI adoption, summarized in Appendix Table LABEL:tab:aiadoption. Prior research has predominantly focused on the antecedents and consequences of adopting traditional, non-generative AI technologies such as automated guided vehicles, computer vision, and natural language processing (McElheran et al. 2024, Chen et al. 2024, Alekseeva et al. 2020). In exploring the antecedents, studies have largely adopted a behavioral perspective to investigate how technological, organizational, and environmental factors influence AI adoption decisions (Jo and Bang 2023, Prasad Agrawal 2023). These studies primarily employ qualitative methods, including surveys and interviews, to deliberate on users’ perceptions of AI (e.g., perceived ease of use and perceived usefulness) and how these perceptions affect adoption incentives (Gupta et al. 2024, Choudhury and Shamszare 2023). Regarding the consequences, it is well established that although AI adoption produces economic benefits, these benefits may be unevenly distributed. For instance, Agrawal et al. (2024) suggests that unequal access to and capabilities in utilizing non-generative AI technologies can exacerbate disparities among firms in terms of organizational performance and market dynamics. This issue is particularly concerning, as it may lead to a widening gap between firms able to harness the benefits of AI and those that cannot, resulting in significant disparities in innovation, productivity, and market competitiveness (Gans 2023).
Our study advances beyond this literature by developing a unified model framework that simultaneously addresses both antecedents and consequences of AI adoption. While prior research has made valuable contributions, the inherently dynamic and intangible nature of user perceptions has posed significant methodological challenges. Traditional approaches struggle to directly measure the underlying utility driving adoption decisions or capture the dynamic nature of the learning process. To address these limitations, we develop a customized Bayesian learning model that formally characterizes the dynamics of generative AI adoption. This structural approach complements existing behavioral studies by moving beyond static snapshots to capture the iterative nature of technology adoption and learning.
2.3 Learning and Decision-Making under Uncertainty
When confronted with ex ante information uncertainty or incompleteness, decision-makers often rely on learning processes to resolve uncertainty and make more informed decisions. Structural Bayesian learning models have thus been proposed to characterize these dynamic learning processes (Wang and Yang 2024, Hu et al. 2022, Erdem and Keane 1996), and our model follows this stream of literature.
These models have been most commonly applied in the context of product attribute assessment, particularly in evaluating product quality (Erdem and Keane 1996). Consumers iteratively update their beliefs regarding product attributes by utilizing various information signals, including direct experience, advertising exposure, and online reviews (Lin et al. 2015, Erdem et al. 2008). The versatility of these models extends to more complex domains, such as pharmaceutical efficacy evaluation (Ching and Lim 2020), consumption of addictive products (Chen and Rao 2020), and financial investment decisions (Zhou 2021). Moreover, within the field of information systems, Bayesian learning models have been adapted to investigate diverse user behaviors, encompassing phenomena such as opinion polarization (Lu et al. 2022), application usage patterns (Zheng et al. 2020, Zhang et al. 2020), rating behaviors (Ho et al. 2017), ideation processes (Huang et al. 2014), and content generation dynamics (Ghose and Han 2011).
Our research context—individuals learning about generative AI through repeated interactions that are subject to randomness—presents an ideal setting for applying structural learning models. The inherent variability in AI system outputs creates uncertainty that users must resolve through experience, making the Bayesian learning framework particularly appropriate. To our knowledge, this study is among the first to utilize a learning model to characterize heterogeneous AI adoption patterns and to quantify the learning process as a mechanism contributing to the AI divide. Furthermore, given the ongoing evolution of AI technologies, our proposed framework offers potential for longitudinal monitoring of dynamic adoption trends.
3 Data, Variable, and Exploratory Analysis
This section describes the data source and variables for the empirical testing. We conduct reduced-form analyses to explore heterogeneous adoption and usage patterns across diverse demographic cohorts. Subsequently, we underscore the imperative for developing a learning model to formally examine underlying mechanisms, based on the exploratory findings.
3.1 Data Source
This study uses a comprehensive dataset comprising detailed daily clickstream records from users, provided by a prominent, anonymized data collection firm. Participants install a clickstream collection application on their devices, which captures a record in the database for each URL visited, including the user ID, access timestamps, and the specific URL. Complementing these online clickstream activities, the firm collects various demographic attributes, such as gender and age, facilitating social-economic analyses. Furthermore, the URLs enable direct extraction of search keywords entered into search engines or specific websites accessed by users.
To elucidate the adoption dynamics of ChatGPT, we analyze a dataset spanning from November 30, 2022 (ChatGPT’s release) to May 31, 2023. The six-month duration provides sufficient time window to observe meaningful patterns in user behavior, allowing us to track how individuals discover, experiment with, and potentially integrate ChatGPT into their routines. Our sample comprises 8,306 users who were active during the observation window and for whom full demographic profiles are available. Within this cohort, 1,102 individuals have interacted with ChatGPT at least once during the observation period.
3.2 Variable Construction
With the aim of investigating the existence of a digital divide, we seek to link individual differences in ChatGPT usage with their socio-demographic and socio-economic attributes. Specifically, we examine how factors such as educational attainment, age, race, gender, and occupational sector influence the AI adoption process—factors consistently highlighted in the literature on the digital divide. For instance, educational attainment has been identified as one of the most robust predictors of digital disparities (Hidalgo et al. 2020, Hsieh et al. 2011, 2008). Digital literacy has also been strongly associated with age, with younger cohorts typically demonstrating higher proficiency (Wang et al. 2024, Jackson et al. 2001). Numerous studies have further identified significant impacts of racial/ethnic variations (Boeing 2020, Fairlie 2004) and gender disparities (Van Dijk and Hacker 2003, Jackson et al. 2001) on the digital divide. Moreover, the occupational and industrial sectors play a crucial role in shaping digital skills and usage patterns (Hidalgo et al. 2020, Venkatesh et al. 2014). To operationalize these factors, we employ a series of variables, whose descriptions and summary statistics are provided in Table 1.
In addition to demographic attributes, we introduce two individual-level variables to characterize general adoption and usage patterns. Specifically, we define $GPTUsed$ to denote whether the individual has first-hand experience with ChatGPT during the observation window. Moreover, for individuals who used ChatGPT, we construct $GPTFrequency$ to measure the frequency of interactions with ChatGPT. Summary statistics for these two variables are presented in Table 1.
Table 1: Summary Statistics (Exploration)
Variable Description N Mean Std Max Min Median
GPTUsed A binary indicator of whether the user uses ChatGPT at least once. 8,306 0.14 0.34 1 0 0
GPTFrequency Average number of ChatGPT access per day since the first usage of ChatGPT. 8,306 0.17 1.27 63.13 0 0
HighEdu A binary indicator of whether the user receives education after high school. 8,306 0.46 0.50 1 0 0
Age The user’s age in integers. 8,306 45.94 14.71 98 18 44
Male A binary indicator of whether the user is male or not. 8,306 0.39 0.49 1 0 0
White A binary indicator of whether the user is a white person. 8,306 0.74 0.44 1 0 1
IT A binary indicator of whether the user works in IT-related industry. 8,306 0.04 0.20 1 0 0
Note: For users that do not use ChatGPT at all, $GPTFrequency$ is filled with zero for summary statistics.
3.3 Exploratory Analyses
Using the constructed variables, we provide exploratory evidence of the potential digital divide. First, the differences in demographic distributions between the overall sample and the ChatGPT user subset are illustrated in Figure 1. The figures suggest that younger, male, non-white, higher-educated individuals, as well as those employed in the IT industry, are more likely to get initial exposure to ChatGPT.
Figure 1: Demographic distributions for all users and active ChatGPT users.
Given that usage frequency is measurable only among users who have accessed ChatGPT, we apply a Heckman two-stage selection model to correct for potential selection bias. In the first stage, we use a probit model to capture the heterogeneity in users’ initial access to ChatGPT based on demographic characteristics. In the second stage, conditional on individuals having used GPT at least once during the observation window, we estimate the heterogeneity in their average daily access following initial exposure. We employ the two-step estimation procedure, with results presented in Table 2.
Overall, the findings indicate that younger, male, non-white, more educated individuals, and those employed in the IT industry are more likely to use ChatGPT and engage with it more frequently. These findings, with the exception of that related to race, are consistent with the previous literature on the digital divide, which has consistently identified that younger individuals (Wang et al. 2024, Hidalgo et al. 2020), males (Van Dijk and Hacker 2003), individuals with higher educational attainment (Hsieh et al. 2011, 2008), and those with more technology-related experience (Hargittai 2003) are less likely to experience digital divides in the adoption and use of ICT. However, our findings regarding race are counterintuitive, as prior literature suggests that white individuals are more likely to adopt and have access to ICT (Boeing 2020, Fairlie 2004). This inconsistency may be explained by the possibility that non-white users face language barriers prior to using ChatGPT, which might actually drive adoption as a means to overcome language-related challenges.
Table 2: Estimation Results (Heckman)
(1) (2)
Variables GPTUsed GPTFrequency
HighEdu 0.366∗∗∗ 5.638∗∗
(0.0376) (2.468)
Age --0.0241∗∗∗ --0.3734∗∗
(0.00139) (0.159)
Male 0.517∗∗∗ 7.664∗∗
(0.0368) (3.368)
White --0.0923∗∗ --1.687
(0.0403) (0.869)
IT 0.420∗∗∗ 6.311∗∗
(0.0750) (2.852)
Constant --0.429∗∗∗ --18.09∗∗
(0.0640) (8.552)
Observations 8,306 8,306
Note: Standard errors in parentheses; ∗∗∗ p$<$0.01, ∗∗ p$<$0.05, p$<$0.1.
3.4 Variable Construction for the Bayesian Learning Model
The above descriptive analyses, however, are insufficient to disentangle utility heterogeneities and learning ability heterogeneities as potential mechanisms underlying the observed adoption patterns. To move beyond static analyses and fully reveal the underlying dynamics, it is necessary to develop a structural model that identifies the learning process while simultaneously accounting for utility differences.
To achieve this objective, we detail the construction process of the variables employed in the Bayesian learning model. First, among all individuals, we select those who have used ChatGPT at least twice to capture the learning process (Ho et al. 2017). We aggregate the clickstream data on a daily basis, resulting in 183 potential observations per individual, each representing one day’s usage. Since we focus on the learning process of ChatGPT, it is appropriate to exclude a user’s records prior to their first exposure to ChatGPT, whether through reading related news or using it directly. Following these aforementioned pre-processing steps, we construct a dataset consisting of 111,371 observations for 1,102 users.
Table 3: Summary Statistics (Main Model)
Statistic Variable N Mean St. Dev. Max Min Median
Observation-level Variables
Total Activities(Search + ChatGPT) $W_{itd}$ 111,371 18.40 43.18 2,658 0 5
ChatGPT Times $W_{its}$ 111,371 1.21 11.09 2,651 0 0
Search Times $W_{itd}-W_{its}$ 111,371 17.19 41.19 1,297 0 4
News Exposure $W_{itn}$ 111,371 0.01 0.44 130 0 0
Individual-level Variables
High-level Education $HighEdu_{i}$ 1,102 0.57 0.49 1 0 1
Age $Age_{i}$ 1,102 39.11 13.62 88 18 38
Male $Male_{i}$ 1,102 0.60 0.48 1 0 1
White $White_{i}$ 1,102 0.66 0.47 1 0 1
IT $IT_{i}$ 1,102 0.09 0.29 1 0 0
Each time an individual $i$ accesses online resources, they may use ChatGPT or traditional search engines [^3] Here, for traditional search engines, we focus on Google and Microsoft Bing, as they account for the majority of the search engine market share.. We define $W_{itd}$ as the total number of activities in a day for a user, and $W_{its}$ represents the daily usage of ChatGPT. The remainder, $(W_{itd}-W_{its})$, corresponds to the usage of traditional search engines.
Given that ChatGPT represents a novel technology, users may rely on their prior experiences with the tool to evaluate its effectiveness and make informed decisions regarding its future use (Ma et al. 2024, Russo 2024). Consequently, ChatGPT usage (measured by $W_{its}$) not only serves as an outcome of user choice but also becomes an information source for the learning process. To account for other online information sources that may influence individuals’ beliefs (Wang and Yang 2024), we also generate a variable $W_{itn}$, which quantifies the number of times a user clicks on ChatGPT-related news articles within a given week [^4] The news outlets used were retrieved from the following website as of April 2024: https://pressgazette.co.uk/media-audience-and-business-data/media_metrics/most-popular-websites-news-us-monthly-3/. The specific processes through which users utilize these variables to revise their beliefs are detailed in Section 4.2. Consistent with the exploratory analysis, we also include $HighEdu_{i}$, $Age_{i}$, $Male_{i}$, $White_{i}$, and $IT_{i}$ as the source of digital divide in the Bayesian learning model. Table 3 presents summary statistics for all these variables.
4 Model
Motivated by our theoretical inquiry and exploratory analyses, we develop a Bayesian learning model to investigate the dynamics of individual adoption behaviors. Whenever a user would like to access resources from the Internet, she must decide whether to start the process with ChatGPT or traditional search engines. This decision is based on the perceived utility derived from each option, which evolves through a learning process influenced by the user’s prior experiences. Consequently, the final choice is determined through utility maximization. The utility function, learning process, and decision-making procedure will be elaborated upon in the following subsections.
4.1 Utility Function
When user $i$ has a general query requiring online resources at time $t$, they may either utilize a traditional search engine or opt for the newly launched ChatGPT, which is trained on a vast corpus of existing information (Xu et al. 2023). To rationalize the decision-making process between adopting ChatGPT versus the traditional search engine (considered as the outside option in this technology adoption scenario), we denote $u_{it1}$ as the utility that user $i$ derives from choosing ChatGPT at time $t$, and $u_{it0}$ as the utility derived from choosing the outside option.
Given that traditional search engines have been active for a long period and users are already familiar with this option, we assume that its utility $u_{it0}$ can be decomposed into a constant term $c$ and a stochastic error term $\varepsilon_{it0}$, in line with the Bayesian learning literature (Zheng et al. 2020):
$u_{it0}=c+\varepsilon_{it0}$ (1)
Similarly, we decompose the utility of choosing ChatGPT into a representative utility term, which can be explained by observable factors, and a stochastic error term $\varepsilon_{it1}$.
$u_{it1}=v_{i}+\varepsilon_{it1}$ (2)
As introduced in RQ2, users with different demographic backgrounds may derive varying levels of utility from using ChatGPT. Therefore, we model each individual’s demographic attributes as contributing linearly to her representative utility, allowing us to characterize the potential utility divide.
$v_{i}=$ $\alpha_{0i}+\alpha_{1}\cdot HighEdu_{i}+\alpha_{2}\cdot Age_{i}+\alpha_{3}\cdot Male_{i}+\alpha_{4}\cdot White_{i}+\alpha_{5}\cdot IT_{i}$ (3)
Note that in other discrete choice settings (Train 2009), only differences in utility are relevant; therefore, only one constant term is needed in the utility functions of the two alternatives to capture the baseline relative differences between ChatGPT and the traditional search engine. However, in our learning model, where we normalize the distribution of individuals’ initial beliefs, both constants can be identified, as discussed in Section 5. Additionally, the parameters $(\alpha_{1},\ldots,\alpha_{5})$ also capture differences in relative utility instead of absolute utility. For instance, a positive $\alpha_{1}$ would indicate that users with higher education tend to derive greater relative gains from using ChatGPT compared to traditional search engines.
However, unlike traditional search engines, ChatGPT has been launched only recently, and individuals may not initially be fully certain about ChatGPT’s utility $u_{it1}$. Therefore, before a user $i$ decides whether to use ChatGPT to fulfill her query, she first forms an expectation of the utility based on the information available up to period $t$ (hereinafter referred to as “perceived utility”), and then choose the alternative that maximizes her expected utility. The perception formation process follows the Bayesian learning framework (Wang and Yang 2024), and will be elucidated in Section 4.2.
Regarding the stochastic error terms in both utility functions ($\varepsilon_{it0}$ and $\varepsilon_{it1}$), they represent utility shocks that influence individuals’ decisions between the two alternatives but are unobserved by researchers. Consistent with the discrete choice literature (Train 2009), we assume that these error terms follow a Type-I extreme value distribution, allowing the expected utility to be linked to the choice probability in a logit functional form.
4.2 Individuals’ Learning Process
As outlined in the previous section, individuals’ choices of tools are influenced by their perceived utility of each option prior to making a decision. This also aligns with the technology adoption literature, which emphasizes the importance of perception in the adoption of newly launched technologies (Ma et al. 2024, Venkatesh et al. 2003). Given that stochastic shocks ($\varepsilon_{it0}$ and $\varepsilon_{it1}$) affecting consumers’ decisions are known to consumers by definition (Zheng et al. 2020, Train 2009), the perceived utility can be represented as follows.
$E[u_{it0}|I_{it}]=E[c|I_{it}]+\varepsilon_{it0}$ (4)
$E[u_{it1}|I_{it}]=E[v_{i}|I_{it}]+\varepsilon_{it1}$ (5)
Here $I_{it}$ represents all the information available to the user $i$ until period $t$ that can help them form perceptions of the utility of the two options for themselves.
Since traditional search engines have been available for an extended period, individuals are already familiar with them, and their beliefs are assumed to have stabilized prior to the observation period. Consequently, similar to other Bayesian learning models (Zheng et al. 2020), the perceived representative utility of the outside option $E[c|I_{it}]$ is assumed to be constant for consumers a priori and is not influenced by new information.
$E[u_{it0}|I_{it}]=E[c|I_{it}]+\varepsilon_{it0}=c+\varepsilon_{it0}$ (6)
In contrast, the utility of the newly introduced ChatGPT is less certain to users, resulting in a dynamic update of their belief $E[u_{it1}|I_{it}]$ over time as they gather more information and experience with the tool. Following the Bayesian learning literature (Wang and Yang 2024, Erdem and Keane 1996), we assume that prior to any direct experience with ChatGPT, individuals possess some prior information regarding the representative utility of using ChatGPT, which is modeled as following a normal distribution $N(v_{0},\sigma_{0}^{2})$. The individual perceives that the actual representative utility $v_{i}$ is drawn from this distribution, where the mean of this distribution represents the mean perception $v_{0}$. $v_{0}$ may differ from the actual value $v_{i}$ to capture potential biases in perception. The variance $\sigma_{0}^{2}$, on the other hand, reflects the level of uncertainty in individuals’ perceptions: all else being equal, a larger variance indicates greater uncertainty about their expected representative utility, making individuals more influenced by incoming information. Therefore, at period 00, the individual’s mean expectation $E[v_{i}|I_{i0}]$ and the variance of such mean expectation $\sigma_{i0}^{2}$ can be expressed as:
$E[v_{i}|I_{i0}]=v_{0}$
$\sigma_{i0}^{2}=E((v_{i}-v_{0})^{2}|I_{i0})=\sigma_{0}^{2}$ (7)
When users actually use ChatGPT, they gain first-hand experience regarding the utility derived from ChatGPT, thereby updating their beliefs about $v_{i}$ and the associated uncertainty. Consistent with the Bayesian learning framework, we denote the information obtained from such experiences as a signal $s_{itk}$, where individual $i$ uses ChatGPT at the $k$-th time during period $t$ (Erdem and Keane 1996). These signals are informative but inherently noisy, due to either variability in the actual representative utility or users’ subjective perception processes (Zheng et al. 2020). To maintain a conjugate update process (van de Schoot et al. 2021, DeGroot 2005), we leverage the properties of normal distributions and assume that the signal follows a normal distribution (Wang and Yang 2024):
$s_{itk}\sim N(v_{i},\sigma_{s_{i}}^{2})$ (8)
Consistent with prior Bayesian learning studies and the concept of rational expectations, the mean of the signal is assumed to equal the true representative utility (Zhang et al. 2020, Huang et al. 2014). Regarding the signal variance, similar to any Bayesian updating process, it influences the relative weight given to prior beliefs versus the incoming signal when constructing the posterior belief, thereby reflecting individuals’ learning speed in converging towards the actual representative utility (Ho et al. 2017, Erdem and Keane 1996). As introduced by our RQ1 (learning divide), individuals with different backgrounds may exhibit heterogeneous learning abilities. To explore such potential disparities, we link observable attributes to the variance as specified in Equation 9. Specifically, within this equation, we apply an exponential transformation to a linear combination of attributes to ensure that the variance $\sigma_{s_{i}}^{2}$ remains positive and covers the entire positive domain.
$\sigma_{s_{i}}^{2}=\exp(\gamma_{0i}+\gamma_{1}\cdot HighEdu_{i}+\gamma_{2}\cdot Age_{i}+\gamma_{3}\cdot Male_{i}+\gamma_{4}\cdot White_{i}+\gamma_{5}\cdot IT_{i})$ (9)
Furthermore, to control for other information sources that might influence individuals’ perceptions, we incorporate news article access based on clickstream data as an additional source of signals. Specifically, at period $t$, when user $i$ clicks on the $j$-th news article that mentions ChatGPT, we model this signal as being drawn from a normal distribution $n_{itj}\sim N(v_{i},\sigma_{n}^{2})$.
We now further illustrate the perception updating process based on signals. To improve estimation efficiency, we aggregate data and analyze users’ learning processes on a daily basis (Zhou et al. 2022). At the beginning of day $(t-1)$, an individual’s perceived representative utility of ChatGPT is characterized by a normal distribution with mean $E[v_{i}|I_{i(t-1)}]$ and variance $\sigma_{i(t-1)}^{2}$. Throughout day $(t-1)$, user $i$ interacts with ChatGPT for $W_{i(t-1)s}$ times and accesses $W_{i(t-1)n}$ related news articles. Using these signals, the user updates their belief according to the Bayesian rule (DeGroot 2005, Erdem and Keane 1996), forming a new belief at the end of the day that the actual representative utility is drawn from the distribution $N(E[v_{i}|I_{it}],\sigma_{it}^{2})$ where
$E[v_{i}|I_{it}]$ $=\frac{\frac{E[v_{i}|I_{i(t-1)}]}{\sigma_{i(t-1)}^{2}}+\sum_{k=1}^{W_{i(t-1)s}}\frac{s_{i(t-1)j}}{\sigma_{s_{i}}^{2}}+\sum_{j=1}^{W_{i(t-1)n}}\frac{n_{i(t-1)j}}{\sigma_{n}^{2}}}{{\frac{1}{\sigma_{i(t-1)}^{2}}+\frac{W_{i(t-1)s}}{\sigma_{s_{i}}^{2}}}+\frac{W_{i(t-1)n}}{\sigma_{n}^{2}}}$ (10)
$\sigma_{it}^{2}$ $=\frac{1}{\frac{1}{\sigma_{i(t-1)}^{2}}+\frac{W_{i(t-1)s}}{\sigma_{s_{i}}^{2}}+\frac{W_{i(t-1)n}}{\sigma_{n}^{2}}}$ (11)
As additional signals are received, the perceived representative utility $E[v_{i}|I_{it}]$ gradually converges towards the true representative utility $v_{i1}$, albeit with some noise, while the associated uncertainty $\sigma_{it}^{2}$ progressively diminishes towards zero.
Finally, according to the Bayesian updating equation, a larger signal variance implies a slower updating process, as the weight assigned to the noisy but unbiased signal is inversely related to the signal variance (Zheng et al. 2020). For instance, a negative $\gamma_{1}$ would imply that, all else being equal, individuals with higher education are able to converge their perceptions $E[v_{i}|I_{it}]$ to the actual value $v_{i}$ more rapidly after receiving the same number of signals compared to those with lower education.
4.3 Unobserved Heterogeneity and Likelihood Function
In this binary choice scenario, each individual selects the option that maximizes their expected utility by comparing $E[u_{it1}|I_{it}]$ and $E[u_{it0}|I_{it}]$. Assuming that the unobserved utility shock follows a Type-1 extreme value distribution, the difference between these utilities is characterized by a logistic distribution (Train 2009). This assumption enables a closed-form expression for the choice probability of selecting ChatGPT, based on the representative perceived utilities ($E[v_{i}|I_{it}]$ and $c$), using the logit model:
$Pr(d_{itw}=1)$ $=\frac{\exp(E[v_{i}|I_{it}])}{\exp(E[v_{i}|I_{it}])+\exp(c)}$ (12)
Here, $d_{itw}$ represents whether individual $i$ chooses ChatGPT at the $w$-th decision point within period $t$. Given that there are $W_{itd}$ decision opportunities in period $t$, the likelihood of observing the individual choosing ChatGPT for $W_{its}$ occasions and opting for search engines for the remaining $(W_{itd}-W_{its})$ occasions can be expressed as:
$l_{it}(W_{itd},W_{its}\mid\bm{\theta})$ $=\left(\frac{\exp(E[v_{i}|I_{it}])}{\exp(E[v_{i}|I_{it}])+\exp(c)}\right)^{W_{its}}\quad\times\left(\frac{\exp(c)}{\exp(E[v_{i}|I_{it}])+\exp(c)}\right)^{W_{itd}-W_{its}}$ (13)
where $\bm{\theta}$ represents the parameter vector governing the individual’s utility function or learning process: $\bm{\theta}=[c,\alpha_{0i},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\gamma_{0i},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5},\sigma_{n}^{2}]$.
Throughout the model setup, heterogeneity in both the utility function and the signal function is captured using observable variables and is assumed to be independent, conditional on these observables. We now relax this assumption by employing a latent class approach (Kamakura and Russell 1989). Specifically, each individual is assumed to have a probability of belonging to a particular latent class, which introduces two class-specific constants related to the utility function and the signal function, respectively. This approach effectively captures and accounts for heterogeneity that is not reflected in the observable variables. Additionally, it allows for unobserved dependencies. For instance, a negative correlation between utility gain and learning speed implies that, all else being equal, individuals who learn more quickly (i.e., have lower variance) also tend to derive relatively higher utility gains from using ChatGPT.
In practice, for latent class 1, we assign the probability of individuals belonging to this class as $Pr(Class_{i}=1)=\frac{1}{1+\exp(\lambda)}$, with associated parameters $\alpha_{0i}=\alpha_{0}$ and $\gamma_{0i}=\gamma_{0}$. For latent class 2, the probability is given by $Pr(Class_{i}=2)=\frac{\exp(\lambda)}{1+\exp(\lambda)}$, with associated parameters $\alpha_{0i}=\alpha_{0}+\Delta\alpha_{0}$ and $\gamma_{0i}=\gamma_{0}+\Delta\gamma_{0}$.
With the incorporation of the latent class structure, the likelihood function can be rewritten accordingly to account for the probabilistic classification of individuals into different latent classes.
$l_{it}(W_{itd},W_{its}\mid\bm{\theta^{\prime}})=$ $\Pr(Class_{i}=1)*l_{it}(W_{itd},W_{its}|\bm{\theta^{\prime}},Class_{i}=1)+$
$\Pr(Class_{i}=2)*l_{it}(W_{itd},W_{its}|\bm{\theta^{\prime}},Class_{i}=2)$ (14)
where $\bm{\theta^{\prime}}$ includes $[c,\alpha_{0},\Delta\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\gamma_{0},\Delta\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5},\sigma_{n}^{2},\lambda]$.
5 Estimation and Identification
5.1 Model Estimation
Based on the likelihood for each (user, period) pair, the total likelihood function can be articulated by considering $I$ users, each observed over $T_{i}$ periods:
$L(\bm{\theta^{\prime}})=\prod_{i=1}^{I}\prod_{t=1}^{T_{i}}l_{it}(W_{itd},W_{its}\mid\bm{\theta^{\prime}})$ (15)
A common approach to estimate the parameters of interest $\bm{\theta^{\prime}}$ is to find those values that maximize the empirical likelihood (Train 2009). However, this approach is not straightforward for the Bayesian learning model given the available data: we can only observe each individual’s attributes such as $HighEdu_{i}$, $Age_{i}$, $Male_{i}$, $White_{i}$, $IT_{i}$, and their usage of ChatGPT ($W_{its}$) and traditional search engines ($W_{itd}-W_{its}$) during each period. We aim to derive the unknown $\bm{\theta^{\prime}}$, but we cannot directly observe users’ perceived representative utility for ChatGPT ($E[v_{i}|I_{it}]$) at each period or the stochastic error terms ($\varepsilon_{it0}$, $\varepsilon_{it1}$). Similar to the mixed logit model, slight modifications can render a fully closed-form expression of the likelihood unfeasible (Train 2009). Although there is a direct mathematical relationship between perceived utility and choice probability under the logit model, deriving the parameters of interest from perceived utility lacks a closed-form solution, as we cannot observe the initial belief distribution $N(v_{0},\sigma_{0}^{2})$ or the exact signals ($s_{itk}$ and $n_{itk}$) (Erdem and Keane 1996).
To address this challenge, we adopt established practices in Bayesian learning with unobservable signals and employ a maximum simulated likelihood estimation method (Wang and Yang 2024, Erdem and Keane 1996). Specifically, signals are drawn from their respective distributions, governed by the current parameter estimates, for each simulation. These signals are then used in the Bayesian updating process to derive perceived utility and subsequently compute the simulated likelihood. After conducting 100 simulations, the average of the simulated likelihoods is used to approximate the likelihood with all the other unknown components, except for $\bm{\theta^{\prime}}$, integrated out. This average likelihood then serves as the objective function for optimization, enabling us to estimate the parameters that maximize the simulated likelihood.
5.2 Identification
In the Bayesian learning process, users learn the true representative utility from signals and gradually update their mean belief from $v_{0}$ to $v_{i}$. Following prior literature (Wang and Yang 2024, Ching et al. 2013), $v_{0}$ and $v_{i}$ cannot be simultaneously identified, and therefore $v_{0}$ is normalized to zero to ensure model identifiability. With $v_{0}$ held constant, the outside option’s representative utility $c$ is identified based on users’ initial choice probabilities between ChatGPT and traditional search engines. A decrease in $c$ corresponds to a relative increase in the net utility from using ChatGPT, thereby increasing the likelihood of ChatGPT adoption in the absence of other information.
Leveraging $c$, the representative utility of ChatGPT $v_{i}$ is identified based on users’ steady-state choice probabilities. The utility parameters ($\alpha_{0i},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5}$) are then derived based on the heterogeneous steady-state choice probabilities among individuals with different backgrounds. For example, an increase in $\alpha_{1}$ implies a greater impact of education on representative utility, making higher-educated users more likely to choose ChatGPT over traditional search engines.
Furthermore, the variance of the signals ($\sigma_{is}^{2}$ and $\sigma_{n}^{2}$) is identified through the variation in users’ belief updates after receiving different signals. As with the identification of the mean value, we cannot simultaneously identify the prior variance and the signal variance. Thus, the prior variance $\sigma_{0}^{2}$ is fixed at $\exp(4)$ to reflect substantial uncertainty that individuals face when they have no prior information about ChatGPT. The core findings are robust to different choices of prior variance values. As individuals receive varying numbers of news signals over time, differences in how their beliefs evolve provide information about $\sigma_{n}^{2}$. Similarly, the signal parameters ($\gamma_{0i},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5}$) are identified based on the variation in learning speeds across individuals with different backgrounds.
Finally, since individuals exhibit different dynamics depending on their latent class (Zheng et al. 2020), the parameter $\lambda$ that governs the class probability, along with the associated class-specific parameters ($\alpha_{0}$, $\Delta\alpha_{0}$, $\gamma_{0}$, and $\Delta\gamma_{0}$), can be identified. Prior to estimating the model with empirical data, a Monte Carlo simulation is conducted to ensure that all parameters can be accurately identified within the proposed model structure and estimation procedure.
6 Results
Following the estimation procedure described earlier, the results are summarized in Table 4. Firstly, in response to RQ2, we document the existence of a utility divide in real-world contexts, demonstrating that different demographic groups derive varying levels of utility from ChatGPT compared to traditional search engines. Specifically, with regard to age and gender, our analysis reveals that younger users and males derive significantly greater benefit from ChatGPT, reporting higher utility levels ($\alpha_{2}=-0.021$ and $\alpha_{3}=0.562$). These findings are consistent with the extant literature on the digital divide, which has repeatedly shown that younger and male users tend to exhibit higher rates of ICT adoption and usage (Yang and Zhang 2023, Hargittai 2002, Jackson et al. 2001). Regarding occupational differences, our model suggests that individuals employed in IT-related sectors experience enhanced utility from ChatGPT ($\alpha_{5}=0.507$), likely due to their higher levels of digital literacy. This observation aligns with previous studies on the digital divide, which have emphasized that individuals with more advanced computer skills are better positioned to utilize ICT effectively (Hargittai 2003, Van Dijk and Hacker 2003). Moreover, recent research on AI adoption, such as that by Alekseeva et al. (2020), has identified a positive relationship between AI adoption and increased firm growth and investment, particularly in IT-intensive industries.
Table 4: Parameters Estimation
Parameter Description Value Std
$v_{0}$ Prior belief on the mean value of representative utility 0 – (fixed)
log($\sigma_{0}^{2}$) Prior belief on the variance of representative utility 4 – (fixed)
$\lambda$ Latent class probability parameter --0.384 $(0.077)^{***}$
$c$ Mean representative utility of traditional search engines 1.411 $(0.013)^{***}$
$\alpha_{0}$ Intercept of ChatGPT’s representative utility --1.560 $(0.024)^{***}$
$log(\Delta\alpha_{0})$ Difference in ChatGPT’s representative utility between two latent classes 0.976 $(0.005)^{***}$
$\alpha_{1}(HighEdu)$ Coefficient of high education levels on ChatGPT’s representative utility --0.468 $(0.011)^{***}$
$\alpha_{2}(Age)$ Coefficient of ages on ChatGPT’s representative utility --0.021 $(0.001)^{***}$
$\alpha_{3}(Male)$ Coefficient of males on ChatGPT’s representative utility 0.562 $(0.016)^{***}$
$\alpha_{4}(White)$ Coefficient of race (white vs non-white) on ChatGPT’s representative utility --0.208 $(0.011)^{***}$
$\alpha_{5}(IT)$ Coefficient of IT practitioners on ChatGPT’s representative utility 0.507 $(0.021)^{***}$
$\sigma_{n}^{2}$ News signal variance 4.842 $(0.028)^{***}$
$\gamma_{0}$ Intercept of usage signal variance 4.900 $(0.153)^{***}$
$\Delta\gamma_{0}$ Difference in usage signal variance between two latent classes 2.031 $(0.011)^{***}$
$\gamma_{1}(HighEdu)$ Coefficient of high education levels on usage signal variance --0.256 $(0.036)^{***}$
$\gamma_{2}(Age)$ Coefficient of ages on usage signal variance 0.029 $(0.005)^{***}$
$\gamma_{3}(Male)$ Coefficient of males on usage signal variance --0.522 $(0.030)^{***}$
$\gamma_{4}(White)$ Coefficient of white people on usage signal variance --0.481 $(0.129)^{***}$
$\gamma_{5}(IT)$ Coefficient of IT practitioners on usage signal variance --0.456 $(0.045)^{***}$
Note: Standard errors in parentheses; ∗∗∗ p$<$0.01, ∗∗ p$<$0.05, p$<$0.1.
While prior research has investigated adoption behavior in real-world settings and evaluated outcome differences, such as productivity, in laboratory settings, the underlying “unobservable” utility has not been directly inferred from real-world data. Our results contribute to the existing literature by explicitly demonstrating the existence of a utility divide based on demographic attributes.
Moreover, while our findings confirm some of literature’s observations regarding observable behavior, we also present evidence that challenges established notions within the digital divide literature. Most notably, concerning education—traditionally considered one of the most significant predictors of the digital divide (Autor et al. 1998)—we observe that users with lower educational attainment experience substantially higher relative utility from ChatGPT ($\alpha_{1}=-0.468$). This result contradicts prior research suggesting that higher education correlates with improved outcomes in ICT adoption (Scheerder et al. 2017). By disproportionately benefiting users with lower educational attainment, generative AI has the potential to mitigate the outcome divide stemming from educational disparities. Although it is empirically infeasible to test the contributing factors of this phenomenon directly, we provide some intuition and call for further research to gather additional data and study the mechanisms. For example, individuals with different educational backgrounds might possess different needs for online resources, and the needs of less-educated users may align more closely with ChatGPT’s capabilities compared to traditional search engines. Conversely, queries from more educated users may be too complex for direct interaction with ChatGPT, making traditional search engines more suitable. Furthermore, less-educated users may have greater difficulty navigating the results from traditional search engines. Since ChatGPT provides answers in natural language that are easier to understand, less-educated users may derive relatively more benefits from interacting with it, whereas this advantage may be less significant for highly educated individuals.
Furthermore, we also find that white users tend to gain comparatively less utility from ChatGPT ($\alpha_{4}=-0.208$), which contradicts previous literature typically finding that white users have greater access to and derive more benefits from digital technologies (Fairlie 2004). This phenomenon may be attributed to the lower language barriers experienced by white users, leading to relatively diminished marginal utility from ChatGPT compared to traditional search engines.
In addition to the divide in actual utility, we observe a significant deviation of $v_{i}$ from $v_{0}$, where $v_{0}$ is normalized to zero. For example, consider a 25-year-old female, non-white user with low educational attainment, not employed in the IT industry, and belonging to latent class 2. For this individual, the actual representative utility $v_{i}$ is underestimated by $v_{0}-(\alpha_{0}+\Delta\alpha+25\cdot\alpha_{2})=-0.569$. On average, the initial belief regarding ChatGPT’s representative utility is approximately 70% lower than its true representative utility for users in latent class 2. Consequently, without an effective learning process, these users may experience substantial welfare losses in situations where ChatGPT would be more suitable than traditional search engines. Conversely, some individuals overestimate their utility gain from ChatGPT, which may lead them to use ChatGPT when traditional search engines would be more appropriate. Given the pronounced discrepancy between initial beliefs and actual utility, it is crucial for individuals to learn and adjust their perceptions to enable more informed decision-making.
However, the ability to update these beliefs may vary significantly among individuals with different demographic attributes, as highlighted by RQ1 (learning divide). By mathematical definition, the larger the signal variance $\delta_{is}^{2}$, the slower the learning process. Our estimation results substantiate the presence of a significant learning divide. Interestingly, while less-educated and non-white individuals generally benefit more from ChatGPT per use in terms of $v_{i}$, their learning speed is significantly slower, as indicated by the signal variance parameters ($\gamma_{1}=-0.256$ and $\gamma_{4}=-0.481$). Thus, without targeted interventions, lower-educated users may adopt ChatGPT at a slower rate and are more susceptible to extreme belief updates, potentially resulting in suboptimal usage patterns. These heterogeneities help reconcile the apparent contradiction between the utility gains per usage and the observed adoption behavior from the digital divide literature: although less-educated and non-white users exhibit higher utility gains per interaction, their slower learning impedes efficient adoption of ChatGPT. Furthermore, our findings indicate that younger and male users are quicker to perceive the true benefits of ChatGPT ($\gamma_{2}=0.029$ and $\gamma_{3}=-0.522$). Likewise, individuals employed in IT-related industries demonstrate faster learning rates ($\gamma_{5}=-0.456$). Across these dimensions, the learning divide mirrors the utility divide, thereby exacerbating the disparity between different demographic groups.
Finally, it is important to note that our results remain robust when subjected to: (i) alternative value of the prior belief variance (hyperparameter); (ii) the inclusion of different news outlets retrieved in October 2024; and (iii) re-estimating the results after the exclusion of outliers. These robustness checks further reinforce the validity of our findings.
7 Policy Simulation
Our empirical results have revealed the presence of both learning and utility divides. While the impact of the utility divide is straightforward, the effect of the learning divide on user adoption outcomes and welfare is more complex. To further elucidate the consequences of the learning divide, including potential negative impacts and possible mitigation strategies (RQ3), we conduct numerical simulations complemented by policy evaluations.
First, to demonstrate the importance of learning in the adoption process, we simulate the belief sequence of a typical fast learner. Based on model estimates, we consider a 28-year-old [^5] The 25th percentile of age is used to represent younger individuals. male IT practitioner with a high level of education from latent class 2. This individual faces the choice between ChatGPT and traditional search engines 18 times [^6] The mean of daily activities is used to represent typical behavior. a day over 1000 days. The red dots in Figure 3 illustrate the simulated trajectory of his belief over time. We observe that the user stops using ChatGPT and updating his belief after receiving relatively negative signals during the initial interactions. Despite substantial uncertainty about the true utility of ChatGPT, the user finds it unrewarding to continue using it and instead sticks with search engines. Consequently, he misses the opportunity to update his belief and potentially benefit from ChatGPT. In other words, a few negative experiences can sometimes lead to a persistent negative belief, even for fast learners. We refer to this phenomenon as a belief trap, where users are trapped by their negative beliefs, cease to engage with new technology, and struggle to escape without external intervention.
Given the significant implications of the belief trap, we investigate how the learning divide influences the likelihood of becoming trapped. In addition to the fast learner, for comparison, we consider a typical slow learner: a 48-year-old [^7] We use the 75th percentile of age to represent relatively elder individuals. female, non-white, non-IT practitioner with a low level of education. To control for the impact of unobservables, we also assign this individual to latent class 2. We mathematically define “being trapped” as a user whose current probability of using ChatGPT, based on their present beliefs, is below $1\%$ of the optimal probability determined by the true mean utility $v_{i}$ after one year [^8] The relative relationships depicted in the subsequent Figure 3 remain consistent across different time horizons; however, we present the snapshot trapping probability after a one-year period only for illustrative purposes.. Accordingly, we simulate 10,000 trajectories for each user and plot the probability of becoming trapped. The first two error bars in Figure 3 depict the simulated probabilities of being trapped for the fast learner and the slow learner, respectively, along with 95% confidence intervals calculated using 1,000 bootstrap samples. The slow learner exhibits a significantly higher probability of being trapped, likely due to greater uncertainty in the learning process. To control for the possibility that this difference is driven by heterogeneous utilities—since higher utilities facilitate adoption and thus belief updates—we further simulate a scenario where the slow learner’s true utility is set equal to that of the fast learner. As shown in the third bar of Figure 3, even after controlling for utility, slow learners remain more likely to become trapped. This persistent difference in trapping probability underscores the impact of the learning divide on dynamic adoption behaviors and welfare.
Due to the substantial impact of the learning divide, mitigating this divide and associated consequences becomes crucial. Our simulation results suggest that offering free training could be an effective solution. In Figure 3, the blue dots represent the updated trajectory for the fast learner. By providing free training at the outset and ensuring that he uses ChatGPT 200 times before Day 00, we observe that he is no longer trapped; his belief is frequently updated and approaches his true representative utility of ChatGPT, $0.90$. These findings demonstrate that early training can help users overcome the belief trap. Therefore, offering training to slow learners may help mitigate the learning divide. Furthermore, the last two error bars in Figure 3 illustrate the simulated probability of the slow learner being trapped after 100 and 150 times of interactions, respectively. The results indicate that a training program involving 100–150 uses of ChatGPT enables slow learners to achieve trapping probabilities similar to fast learners, suggesting that this approach can effectively mitigate the learning divide in practice.
Figure 2: Belief Change Over Time
Figure 3: Simulated Probability of Being Trapped
8 Conclusion
Inspired by prior literature on the digital divide, our study introduces two novel concepts to better understand the generative AI divide. First, the learning divide, a critical facet of the capability divide, highlights disparities in individuals’ abilities to effectively update their perceived utility of generative AI through repeated interactions and access to information resources. Second, the utility divide, a form of outcome divide, characterizes differences in the utility derived from each usage of ChatGPT among individuals with different demographics. Utilizing a six-month clickstream dataset, we build and estimate a Bayesian learning model to examine how different demographic groups learn about and derive utility from ChatGPT.
Our analyses reveal significant learning and utility divides across demographics. Males, younger individuals, and those with IT backgrounds not only derive higher utility per use but also update their beliefs about GenAI more rapidly. Conversely, individuals from minority racial groups or with lower educational attainment update their beliefs more slowly, even though they may experience higher utility gains per use. Building on these empirical insights, we conduct simulations using the estimated parameters to model belief sequences over time. Notably, we identify a concerning phenomenon—the belief trap—where users who underestimate generative AI’s utility may abandon the technology, foregoing opportunities for welfare enhancement. The simulations further demonstrate how the learning divide leads to significant differences in the probability of being trapped, thereby contributing to broader outcome divides in adoption dynamics. Finally, we employ counterfactual simulations to show that interventions such as training programs can mitigate these disparities by enhancing exposure and improving belief updating processes.
This study makes several contributions to the literature on the digital divide, AI adoption, and Bayesian learning. First, we develop a novel framework for understanding the AI adoption process in contexts characterized by information uncertainty and dynamic learning. Unlike previous research that often focuses on access and initial usage, our model emphasizes the ongoing learning dynamics that shape user interactions with AI technologies over time. Second, we introduce the concepts of the learning divide, utility divide, and belief trap to elucidate the mechanisms behind differential AI adoption and outcomes. Third, we highlight how AI technologies, without proper intervention strategies, risk amplifying existing social inequalities, emphasizing the critical importance of inclusive design and deployment practices. This highlights the importance of considering equity and inclusion in the design and deployment of AI systems.
Beyond theoretical contributions, our findings have practical implications for governments, NGOs, AI providers, and companies. Governments and NGOs should promote users’ capabilities in using AI through educational programs to help underprivileged users overcome information asymmetry. AI platforms, such as Hugging Face, could implement inclusivity-focused ELO rankings, create inclusive benchmarks, and introduce badges to motivate developers towards inclusivity. Companies that provide AI or integrate AI into their products should prioritize inclusive product design and offer training programs for disadvantaged groups. By addressing the utility divide, learning divide, and belief trap, these stakeholders can promote equitable AI adoption and ensure that the benefits of AI technologies are broadly shared across society.
Despite these contributions, our work has limitations that warrant future research. While we illuminate the learning divide and belief trap in individual adoption of generative AI, future research could investigate these phenomena in organizational settings, where factors such as organizational culture and collective learning may influence the AI adoption process. Furthermore, our Bayesian learning model could be applied to other emerging technologies characterized by uncertainty and learning dynamics, such as blockchain (Gao et al. 2021), to determine whether similar divides exist. Finally, future research could explore how disparities in AI adoption affect market competition and economic inequalities. We call for more interdisciplinary research to further understand these issues and to develop strategies that promote equitable AI adoption, ultimately informing policymakers and organizations in fostering inclusive technological progress.